OpenStudy (anonymous):

A wheel has a constant angular acceleration of 3.3 rad/s^(2). During a certain 6.0 s interval, it turns through an angle of 93 rad. Assuming that the wheel started from rest, how long had it been in motion before the start of the 6.0 s interval?

OpenStudy (michele_laino):

here we can write this ODE: \[\frac{{d\omega }}{{dt}} = k\] where \(k=3.3\)

OpenStudy (michele_laino):

let's suppose that the 6 seconds interval starts at time \(t_1\), then the angular speed and the angular displacement, at \(t_1\), are: \[\omega \left( {{t_1}} \right) = k{t_1},\quad \theta \left( {{t_1}} \right) = k\frac{{t_1^2}}{2}\] repsectively. Now, at time \(t_1\) starts the 6 seconds interval, and final time is \(t_2\), namley: \(t_2-t_1=6\) so, integrating twice that ODE, with the initial condition above, we get: \[\Delta \theta = \theta \left( {{t_2}} \right) - \theta \left( {{t_1}} \right) = k{t_2}\left( {\frac{{{t_2}}}{2} + {t_1}} \right)\] where \( \Delta \theta =93\). Remembering that \(t_2=t_1+6\), we can easily compute the value of \(t_1\)

OpenStudy (michele_laino):

oops.. namely*